A survey of 1,000 passengers at KLIA has found that the time taken to
get between 2 gates in the same terminal of the airport is normally
distributed with a mean of 20 minutes and a standard deviation of 6
minutes.
a) Based on this information, what should be the minimum time allowed
for a connection between an arrival and a departure in order to ensure
that the majority of passengers are able to make it?
b) If an arriving plane is late, reducing the available connection time to 30
minutes, what percentage of passengers would you expect to arrive
late at the departure gate?

Answer by ewatrrr(24785)

mean of 20 minutes and a standard deviation of 6 minutes.
b) P(x≥ 30) = P(z ≥ 10/6 = P(z ≥ 1.6667) = 0.0478 0r 4.78%